OpenStudy (anonymous):

A stone thrown vertically upward was noted having a velocity of 18.54 m/s after covering half its maximum height. a. what is the total height reached by the ball? b. with what velocity was it thrown? c. how long will it take for the ball to acquire a velocity of 10.5m/sec on its ways down? Really need :(

OpenStudy (john_es):

Using energy conservation, in the midpoint, a) \[\frac{1}{2}mv^2=mg\frac{h}{2}\Rightarrow h=v^2/g=18.54^2/9.8=35.07\ m\] b) \[v_0=\sqrt{2gh}=26.22\ m/s\] c)\[10.5=gt\Rightarrow t=1.07\ s\]